# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v1

@@ III-B3 Generator Enhancement
Decoding tuning refers to adding additional controls during the generator processing, which can be achieved by adjusting hyperparameters to achieve greater diversity, limiting the output vocabulary in some form, and so on. InferFix [131] balances the diversity and quality of results by adjusting the temperature in decoder. SYNCHROMESH [120] limits the output vocabulary of the decoder by implementing a completion engine to eliminate implementation errors. DeepICL [147] controls the randomness of the generation according to an additional temperature factor.
@@ III-B5 RAG Pipeline Enhancement
Model-based: Self-RAG [126] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different instructions, and evaluates the relevance and level of support of the retrieved text through the Self-Reflection token. Finally, the quality of the final output result is evaluated based on the Critique token. Ren et al. [195] uses ”Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [196] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed.
@@
RepoCoder [197] applies an iterative retrieval-generation pipeline to better utilize the useful information scattered in different files in the code completion tasks. It augments the retrieval query during the i-th iteration with previously generated code and obtain greater results. ITER-RETGEN [198] synergizes retrieval and generation in an iterative manner. The current output of the generator can to some extent reflect the knowledge it still lacks, and the retrieve can retrieve the missing information as contextual information for the next round, which helps to improve the quality of the generated content in the next round.
@@ IV-A1 Question Answering
FID [33] and REALM [31] identify the top k most pertinent article snippets based on the query and forward each snippet along with the question to large language models (LLMs) to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [199] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. Atlas [28] demonstrated that leveraging Retrieval-Augmented Generation (RAG) to assist LLMs in open-domain question answering tasks significantly enhances few-shot learning capabilities. RETRO [34] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [196] observed that using RAG does not invariably benefit Question Answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement.KAPING [200] enhances the quality of LLM responses by incorporating structured knowledge from knowledge graphs into the input via constructed prompts. TOG [201] introduced an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the Knowledge Graph and by expanding the inference path space with beam search. Addressing the long-tail issue. NPM [160] pioneered the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. Self-RAG [126] improves answer quality by learning to discern when to retrieve, assess the retrieved content’s relevance, and evaluate the final generated results using four types of reflective tokens. CL-ReLKT [202] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [203] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [155] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing. UR-QA [204] found that when encountering unseen problems, retrieving QA pairs has a better final effect; When encountering problems that have not been seen before, the retrieve text chunk performs better. Therefore, it is proposed to simultaneously retrieve QA pairs and text chunks, and select the final answer by comparing the calibrated confidences.
@@ IV-A2 Fact Verification
CONCRETE [205] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. However, the vast scope of common sense knowledge and the intricacies of reasoning processes make Commonsense Reasoning a perennially challenging and prominent area of research within the field of Natural Language Processing (NLP).
KG-BART [206] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid large language models (LLMs) in crafting more nuanced and logically coherent sentences. This approach not only improves the models’ generalization capabilities but also significantly bolsters their Commonsense Reasoning proficiency.
@@ IV-A4 Human-Machine Conversation
ConceptFlow [207] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. This method achieves commendable results even with a substantial reduction in model parameters. Cai et al. [208] reimagines the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [209] augments dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [210] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [211] improves the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language. It is a pivotal task in the domain of Natural Language Processing (NLP) and represents a significant objective in the pursuit of Artificial Intelligence (AI), boasting considerable scientific and practical significance.
@@ IV-A7 Summarization
In the realm of Natural Language Processing (NLP), Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. Summarization enables users to quickly grasp the essence of a text, thereby conserving time that would otherwise be spent on reading extensive material. There are two main approaches to Summarization: Extractive and Abstractive.
Extractive Summarization involves the automatic selection and compilation of key phrases directly from the source text. A key phrase succinctly captures the main themes, content, or perspectives of the text and is typically composed of one or several words. The generation of key phrases is instrumental for understanding, categorizing, retrieving, and organizing textual information. It is extensively applied in fields such as search engine optimization, academic research, text summarization, and more. This technique refrains from creating new sentences, instead repurposing segments from the original text.
Abstractive Summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences. This approach can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity.
@@ IV-B RAG for Code
Code-related tasks, including code generation, code summarization, code auto-completion, automatic program repair, etc., are always a focus in the field of software engineering. Retrieval and generation are previously two separate ways to resolve these tasks. For retrieval, similar code snippets can be found through similar Abstract Syntax Trees or text edit distances, while natural languages can be found through sparse retrieval or dense retrieval. For generation, considering that code has a sequential data form, many sequence-to-sequence models are applied, from LSTM, Transformers, to current decoder-only LLMs. Recent research leverages both techniques to achieve better performance, augmenting the generation process with the retrieved results.
@@ IV-B1 Code Generation
Retrieval-based prompt engineering in few-shot learning is one of the most prevalent scenarios of RAG in code generation. Few-shot learning, or say in-context learning, includes training samples in prompts as the input for sequence-to-sequence generative models. Retrieval techniques are adopted to find similar training samples to the test input, so that the prompt can be more informative and related. REDCODER [38] retrieves similar NL descriptions using dense retriever CodeBert [24], then concatenates the NL texts, their paired codes, for downstream generator PLBART [39]. APICoder [129] first train a Bert-based deep retriever to align the embeddings of NL descriptions and API documentation; then, it retrieves relevant API information to build prompt for the generator CODEGEN-MONO [216]. COCOGEN [217] aims at commonsense reasoning, which generates code-based reasoning-graphs with Codex [2] given NL inputs; it adds an evaluation setting of dynamic prompt selection, which actually retrieves relevant examples for prompts. In DocPrompting [40] given an NL intent, the retriver retrieves relevant documentations, then the generator generates codes based on the NL and retrieved documents. It evaluates both sparse retrievers and dense retrievers, and also tries different generators in experiments. CodeT5+ [218] adopts the CodeT5 [105] model for both the retriever and generator, leveraging only the encoder part in the retrieval process. AceCoder [177] fixes the retriever to BM25 [18], and tests several LLM generators for code generation.
@@ IV-B2 Code Summary
In many research works, the retrieval results are processed by additional encoders. Rencos [119] utilizes two different methods to retrieve similar code snippets, syntactic similarity between abstract syntax trees (AST) and cosine similarity between dense representations. For generator, it adopts attention-based LSTM. There are three encoders for the code input and two retrieval results respectively, and the decoder combines three probabilities for final generation. Re2Com [220] and EditSum [138] both retrieve similar code using sparse retrieval BM25 and generate summary using LSTM. They separately encode the input, the retrieved code, and the corresponding summary, then combine the middle representations or probabilities in the decoder. HGNN [221] instead uses code edit distance for retrieval, and substitutes the encoders for codes with hybrid GNN on their Code Property Graphs (CPG) [222]. RACE [141] aims at generating commit message for code difference. It leverages dense retrieval for similar examples and transformer model for generation. It also uses three encoders for the input, the retrieved code difference, and corresponding commit message, then combines the results before feeding into the decoder. BASHEXPLAINER [139] shares the similar idea. Its dense retrieval module is based on CodeBert [24], and for generation, the output representations of the input and the similar code from CodeBert are directly fused for transformer-based decoder.
RAG for in-context learning, which retrieves similar examples and build prompt for generation, also works in code summary. REDCODER [38] works for both code generation and summary, and it replaces the retriever with GraphCodeBert [104] for code summary task. ASAP [185] retrieves similar code with BM25, and generates summary with GPT models.
@@ IV-B3 Code Completion
Code completion can be thought of as the coding equivalent of the “next sentence prediction” task. ReACC [132] follows the few-shot learning paradigm, retrieving similar codes to build prompts for generation. For retrieval, it uses hybrid retriever, which combines sparse and dense retrieval; for generation, it uses CodeGPT-adapted [223]. RepoCoder [197] steps further to perform iterative retrieval and generation to bridge the gap between the retrieval context and the intended completion target. In each iteration, for retrieval, the code query is augmented with previously generated code; for generation, the prompt is formed by combining the newest retrieved codes with the code query. Other than combining retrieval results in prompts, the retrieval-and-edit framework [140] first retrieves similar training examples using dense retrieval, then encodes the input and the retrieved result separately, finally combine them through attention for later LSTM decoder. CoCoMic [224] builds a project context graph for the whole code project, and retrieves top-k neighbors of the input source code. It generates representations of both source code and retrieved contexts, then jointly processes the embeddings to complete the code.

### v2

@@ III-B3 Generator Enhancement
Decoding tuning refers to adding additional controls during the generator processing, which can be achieved by adjusting hyperparameters to achieve greater diversity, limiting the output vocabulary in some form, and so on. InferFix [132] balances the diversity and quality of results by adjusting the temperature in decoder. SYNCHROMESH [121] limits the output vocabulary of the decoder by implementing a completion engine to eliminate implementation errors.
@@ III-B5 RAG Pipeline Enhancement
Model-based: Self-RAG [127] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different instructions, and evaluates the relevance and level of support of the retrieved text through the Self-Reflection token. Finally, the quality of the final output result is evaluated based on the Critique token. Ren et al. [218] uses ”Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [219] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed. Rowen [220] employs a model as a sophisticated multilingual detection system to evaluate the semantic coherence of answers to identical questions posed across various languages. In the event of detected inconsistencies, it decides to retrieve external information, thereby enhancing the reasoning process and rectifying inaccuracies. Conversely, when responses exhibit consistency, the system upholds the initially generated answer, which is derived from internal reasoning. AdaptiveRAG [221] dynamically decides whether to retrieve based on the query complexity by a classifier, which is a smaller LM.
@@
RepoCoder [222] applies an iterative retrieval-generation pipeline to better utilize the useful information scattered in different files in the code completion tasks. It augments the retrieval query during the i-th iteration with previously generated code and obtain greater results. ITER-RETGEN [223] synergizes retrieval and generation in an iterative manner. The current output of the generator can to some extent reflect the knowledge it still lacks, and the retrieve can retrieve the missing information as contextual information for the next round, which helps to improve the quality of the generated content in the next round. SelfMemory [224] employs a retrieval-augmented generator in an iterative manner to create an unlimited memory pool. Following this, a memory selector is used to choose one output, which then serves as the memory for the subsequent generation round.
@@ IV-A1 Question Answering
FID [34] and REALM [32] identify the top k most pertinent article snippets based on the query and forward each snippet along with the question to LLMs to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [225] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. Atlas [29] demonstrated that leveraging RAG to assist LLMs in open-domain question answering tasks significantly enhances few-shot learning capabilities. As shown in Fig. 9, RETRO [35] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [219] observed that using RAG does not invariably benefit Question Answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement. TOG [226] introduced an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the Knowledge Graph and by expanding the inference path space with beam search. NPM [161] pioneered the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. Self-RAG [127] improves answer quality by learning to discern when to retrieve, assess the retrieved content’s relevance, and evaluate the final generated results using four types of reflective tokens. CL-ReLKT [227] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [228] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [154] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing.
Figure 9: Architecture of RETRO [35] model.
UR-QA [229] found that when encountering unseen problems, retrieving QA pairs has a better final effect; When encountering problems that have not been seen before, the retrieve text chunk performs better. Therefore, it is proposed to simultaneously retrieve QA pairs and text chunks, and select the final answer by comparing the calibrated confidences. ARM-RAG [230] demonstrates exceptional performance in solving elementary mathematics problems by retrieving past successful reasoning steps. DISC-LawLLM [231] constructs a supervised fine-tuning dataset through a legal syllogism prompting strategy, enabling the model to receive support from the latest legal information. RAG-end2end [232] conducts simultaneous training of the retriever(DPR) and the generator(BART) to optimize performance for the end-to-end question-answering task and to facilitate domain adaptation. MultiHop-RAG [233] is designed to extract pertinent information from a variety of distinct documents, aggregating this knowledge to equip the generator with the necessary context for producing the definitive answer to the query.
@@ IV-A2 Fact Verification
CONCRETE [234] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [235] proves on LLaMA [236] and Atlas [29] that search augmentation is more beneficial for solving inconsistency problems than increasing model size.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. However, the vast scope of common sense knowledge and the intricacies of reasoning processes make Commonsense Reasoning a perennially challenging and prominent area of research within the field of NLP.
KG-BART [237] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. This approach not only improves the models’ generalization capabilities but also significantly bolsters their Commonsense Reasoning proficiency. Wan et al. [238] constructs the CONFLICTINGQA dataset, comprising contentious questions and conflicting answer documents, to examine which textual features significantly influence LMs’ ability to independently navigate controversial issues. The findings reveal that LMs often neglect the stylistic aspects of text that are typically valued by people.
@@ IV-A4 Human-Machine Conversation
ConceptFlow [239] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. This method achieves commendable results even with a substantial reduction in model parameters. Cai et al. [240] reimagines the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [241] augments dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [242] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [243], PARC [244], and CREA-ICL [245] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [246] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval. If the answer is correct, the retrieved document is added to the original answer as a reference; if the answer lacks reliable references, it guides the LLM to respond to the question anew.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language. It is a pivotal task in the domain of NLP and represents a significant objective in the pursuit of AI, boasting considerable scientific and practical significance.
@@ IV-A7 Summarization
In the realm of NLP, Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. Summarization enables users to quickly grasp the essence of a text, thereby conserving time that would otherwise be spent on reading extensive material. There are two main approaches to Summarization: Extractive and Abstractive.
Extractive Summarization involves the automatic selection and compilation of key phrases directly from the source text. A key phrase succinctly captures the main themes, content, or perspectives of the text and is typically composed of one or several words. The generation of key phrases is instrumental for understanding, categorizing, retrieving, and organizing textual information. It is extensively applied in fields such as search engine optimization, academic research, text summarization, and more. This technique refrains from creating new sentences, instead repurposing segments from the original text. Abstractive Summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences [250, 156, 251, 252]. This approach can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity.
@@
RPRR [251] employs a Retrieve-Plan-Retrieve-Read approach to overcome the limited context window constraints faced by LLMs, utilizing retrieved information to generate high-quality Wikipedia documents for emerging events. RIGHT [252] chooses to use different types of retrievers in different datasets to enhance the generator, which can effectively improve the quality of automatically generated labels in simple deployment.
@@ IV-B RAG for Code
Separate retrieval and generation approaches have historically been employed for code-related tasks. For retrieval, similar code snippets can be identified using Abstract Syntax Trees or text edit distances, while natural languages can be retrieved through sparse or dense retrieval. For generation, sequence-to-sequence models (such as LSTM and Transformers) are employed to generate code or natural language in sequential data form. Recent RAG research combines both retrieval and generation techniques to enhance the overall performance.
@@ IV-B1 Code Generation
Retrieval-based prompt engineering is one of the most prevalent scenarios of RAG in code generation. In-context learning includes training samples in prompts as the input for sequence-to-sequence generative models. Retrieval techniques are adopted to find similar training samples to the test input, so that the prompt can be more informative and related. REDCODER [39] retrieves similar NL descriptions using dense retriever CodeBert [25], then concatenates the NL texts, their paired codes, for downstream generator PLBART [40]. CodeT5Mix [253] adopts the paradigm of RECODER, proposing a model that functions dually as both the retriever and generator. APICoder [130] first train a Bert-based deep retriever to align the embeddings of NL descriptions and API documentation; then, it retrieves relevant API information to build prompt for the generator CODEGEN-MONO [254]. The following work [255] uses the same pipline, renaming the retriever module as APIFinder. COCOGEN [256] aims at commonsense reasoning, which generates code-based reasoning-graphs with Codex [2] given NL inputs; it adds an evaluation setting of dynamic prompt selection, which actually retrieves relevant examples for prompts. In DocPrompting [41] given an NL intent, the retriver retrieves relevant documentations, then the generator generates codes based on the NL and retrieved documents. It evaluates both sparse retrievers and dense retrievers, and also tries different generators in experiments. CodeT5+ [257] adopts the CodeT5 [106] model for both the retriever and generator, leveraging only the encoder part in the retrieval process. AceCoder [193] fixes the retriever to BM25 [19], and tests several LLM generators for code generation. A3CodGen [258] extracts local information, retrieves relevant global functions using embeddings, and incorporates third-party library information to construct the prompt for LLM-based code generation. SKCODER [259] employs BM25 to retrieve relevant code snippets, which are further processed to produce sketch template. The template and the original description are concatenated for final generation. CodeGen4Libs [260] uses RAG for both import statements and codes, employing BM25 as retriever and finetuned CodeT5 as generator. CODEAGENT [261] design agents to search on web, retrieve relevant documentation, generate programs, and test correctness. RRGCode [262] retrieves relevant code snippets using both sparse and dense retrieval, employs a cross-encoder to re-rank the retrieval results, then generates code using the concatenation of the query and the retrieved codes. A recent study [263] shows that retrieval-augmented framework for code suggestions, including code generation and code completion, can improve the performance by a large margin. ARKS [264] steps further upon prompt-based RAG, incorporating iterative RAG to re-formulate queries and update knowledge soup (containing documentation, execution feedback, generated code, etc.) for dense retrieval, improving final generation accuracy.
@@
ToolCoder [266] performs normal code generation, and conducts online search or offline retrieval when encountering special ¡API¿ token. This paradigm makes the model learn to leverage API tools.
@@ IV-B2 Code Summary
In many research works, the retrieval results are processed by additional encoders.
Figure 10: Architecture of Re2Com [142] model.
Re2Com [142] and EditSum [139] both retrieve similar code using sparse retrieval BM25 and generate summary using LSTM. As shown in Fig. 10, they separately encode the input, the retrieved code, and the corresponding summary, then combine the middle representations or probabilities in the decoder. HGNN [267] instead uses code edit distance for retrieval, and substitutes the encoders for codes with hybrid GNN on their Code Property Graphs (CPG) [268]. RACE [143] aims at generating commit message for code difference. It leverages dense retrieval for similar examples and transformer model for generation. It also uses three encoders for the input, the retrieved code difference, and corresponding commit message, then combines the results before feeding into the decoder. BASHEXPLAINER [140] shares the similar idea. Its dense retrieval module is based on CodeBert [25], and for generation, the output representations of the input and the similar code from CodeBert are directly fused for transformer-based decoder. READSUM [269] retrieves similar code using Levenshtein distance, applies code encoder and summary encoder to retrieved pairs, and generates summary using a decoder where a fusion network combines the information.
RAG for in-context learning, which retrieves similar examples and build prompt for generation, also works in code summary. REDCODER [39] works for both code generation and summary, and it replaces the retriever with GraphCodeBert [105] for code summary task. ASAP [208] retrieves similar code with BM25, and generates summary with GPT models. Similar to the paradigm described above, research on pseudocode generation [270] employs the retrieved code as input for generation and subsequently replaces the results with the original input. SCCLLM [271] retrieves relevant code snippets by semantic, syntactic, and lexical-based retrieval, then forms in-context prompts for smart contract comment generation via LLM. UniLog [272] retrieves similar code snippets paired with their log statements to conduct in-context learning for log statement generation.
Logit-based RAG also prevails in code summarization. Rencos [120] utilizes two different methods to retrieve similar code snippets, syntactic similarity between abstract syntax trees (AST) and cosine similarity between dense representations. For generator, it adopts attention-based LSTM. There are three encoder-decoder models for the code input and two retrieval results respectively, and the probabilities are combined for final generation. CoRec [273] generates commit message given code diff and retrieved similar code diff. Multiple LSTM generators handle the input code diff and the retrieved code diff, then adds the probabilities for final generation. kNN-Transformer [274] obtains context vector by feeding code into an encoder-decoder generator, then gets three parts of logits: the first is from searching the vector, the second is from normal Transformer, and the third is the copy mechanism that reserve rare tokens from the input. Tram [275] involves retrieval in both token-level and sentence-level. It encodes source code and corresponding AST into hidden states representations; for token-level, it retrieves similar representations in the database to form next-token prediction logits; for sentence-level, it uses similar code for autoregressive generation logits; it also add the original autoregressive generation logits. CMR-Sum [276] uses encoder-decoder model to generate summaries. It conducts cross attention between representations of retrieved summary and generated summary, and add the logits to the original transformer logits for final distribution.
@@ IV-B3 Code Completion
Code completion can be thought of as the coding equivalent of the “next sentence prediction” task. Early attempts on function completion [277] employs DPR to retrieve relevant template functions using function docstring, then concatenate the information as the input to the code generator BART. ReACC [133] retrieves similar codes to build prompts for generation. For retrieval, it uses hybrid retriever, which combines sparse and dense retrieval; for generation, it uses CodeGPT-adapted [278]. RepoCoder [222] steps further to perform iterative retrieval and generation to bridge the gap between the retrieval context and the intended completion target. In each iteration, for retrieval, the code query is augmented with previously generated code; for generation, the prompt is formed by combining the newest retrieved codes with the code query. Other than combining retrieval results in prompts, the retrieval-and-edit framework [141] first retrieves similar training examples using dense retrieval, then encodes the input and the retrieved result separately, finally combine them through attention for later LSTM decoder. CoCoMic [279] builds a project context graph for the whole code project, and retrieves top-k neighbors of the input source code. It generates representations of both source code and retrieved contexts, then jointly processes the embeddings to complete the code. RepoFusion [280] follows the idea of Fusion-in-Decoder; it employs multiple encoder to input the concatenation of the retrieved repo contexts and the unfinished code, them fuses them and generates the results through a decoder. KNM-LM [281] uses the same model for retrieval encoding and generation, then combines the logits using bayes inference. EDITAS [282] aims at assertion generation. It retrieves similar queries and their assertions, encodes the edit sequence (from the retrieved query to the original query) and the retrieved assertion, then fuses the information for decoder generation. De-Hallucinator [283] first generates code snippets without retrieval, then retrieves relevant API references given generated contents. In the second pass, retrieved API references are combined in prompt for better generation. REPOFUSE [284] uses both rationale context and retrieved similar codes to construct prompt. To fit in the input length limit, it reserves the contexts that are most relevant to the query.